The sum of two numbers is -64. Twice the first number is 1 more than the second. Find the two numbers

Answer by drk(1908)

Let X = first number and Y = second number.
The sum of two numbers is -64 can be translated as
(i) X + Y = -64
Twice the first number is 1 more than the second can be translated as
(ii) 2X = Y + 1.
We can use substitution to solve. Solve the (ii) for Y. We get
(iii) Y = 2X - 1
Substitute (iii) into (i) to get
(iv) X + 2X - 1 = -64
Solve for X.
3X - 1 = -64
3X = -63
X = -21
Find Y.
Y = -43.